The impact of financial leverage on return on equity and earnings per share Consider this case: Rinsemator Group. is considering a project that will require $350,000 in assets The project is expected to produce an EBIT (earnings before interest and taxes) of $55,000 · The project will be financed with 100% equity . There will be 25,000 shares of common equity outstanding · The company faces a tax rate of 30% Using the preceding information, what will be Rinsemator Group.s return on equity (ROE) for this project? 12.10% 11.00% 9.90% 8.80% Rinsemator Group.s earnings per share (EPS) will be if it finances this project with 100% equity Rinsemator Group.s CFO is also considering financing this project with 50% debt and 50% equity. The interest rate on the companys debt will be 13%. Because the company will finance only 50% of the project with equity, it will have only 12,500 shares outstanding. What will be the ROE on this project if the company decides to finance the project with 50% debt and 50% equity? 14.19% 16.13% 12.26% 12.90% O Rinsemator Group. s EPS will be if it finances this project with 50% equity and 50% debt. Typically, using financial leverage will a projects expected ROE
Answer #1

1)Earning after tax : EBIT [1-tax]

            = 55000 [1-.30]

           = 38500

Total asset=total equity = 350000

ROE= net income /Total equity

        = 38500/350000

        = .11 or 11%

correct option is "B"

2)Earning per share =Net income /shares outstanding

               = 38500/ 25000

                = $ 1.54 per share

3)Debt : 350000* .50 =175000

Equity : 350000-175000=175000

Earning before Tax :EBIT -interest

                 = 55000 - [175000*.13 ]

              =55000- 2750

              = 32250

Earning after tax :EBT [1-tax]

            = 32250[1-.30]

            = 22575

ROE =net income /Equity

      =22575 /175000

          = .129 or 12.9%

correct option is "D"

EPS = 22575 / 12500 =$ 1.806

using financial leverage increase a project expected ROE   [from 11% to 12.9%]